OpenStudy (anonymous):

what massof calcium oxide CaO,forms when 25g calcium carbonate CaCO3 decomposes on heating?

OpenStudy (aaronq):

Moles of CaCO3 = m/MW= (25g)/(100g/mol) = 0.25 moles CaCO3 -> CaO + CO2 Coefficients 1 1 1 Moles 0.25 x x Ratio: CaCO3:CaO 1:1 0.25 moles of CaO formed Mass CaO = (MW)(moles)= (56g/mol)(0.25)= 14g